# Archiving Body Movements: Collective Generation of Chinese Calligraphy

### v1

@@ Abstract
Wushu also serves as a virtual platform to review and appreciate the audience-generated works; the audience assumes the role of both writers and readers in the context of the “Book.” Creating (“Writing”) and appreciating (“Reading”) the generated calligraphy becomes a cyclical process within this infinite “Book.” Wushu can stimulate further attention and discussions concerning Chinese characters and calligraphy, such as their visual forms, linguistic meanings, and cultural essence.
@@
However, Wushu has its limitations, challenges, and opportunities that we would like to address in the future. The compendium is an essential component of Wushu’s concept and the archive. The audience finds it intriguing that their collective efforts will form an actual “book.” However, we have not yet figured out the best format to present the archive, except for the real-time showcasing as part of Wushu’s interactive experiences. We currently save the digital pages as image files but are considering an online gallery that updates in real-time and publishes the archive publicly online or a physical book in print that mimics the reading experience. We will further explore and investigate how to present the archive in the future better.
During the exhibition of Wushu, we notice a significant difference in the performance of the English and Chinese caption models. The English caption model, which we acquired from a public online source, generated more accurate and fluent captions than the Chinese caption model, which we obtained from a different source. We attributed this difference not to the lack of AI research on Chinese captioning but to the scarcity and quality of the data sources. The large models that we used for captioning were trained with semantic information from various online sources, but the predominant source of the Chinese data was taobao.com, which had lower quality and diversity than the English data. Although this manuscript and research do not address the data and AI bias issues, we intend to explore them further.
Appendix 01
The optical flow algorithm and partial differential equation (PDE) formula used for the ink diffusion simulation effects:
Where:
  • $\frac{\partial T_{t}}{\partial t}$ is the ink texture changing rate;
  • $\alpha\,\beta$ are the speed parameters controls ink spreading anf absorption respectively;
  • $\nabla^{2}\,\nabla$ are the laplacian and gradient operators for diffusion simulation;
  • $N(x,y,z,t)$ represents Simple3D noise with values in the range [0, 1];
  • $O_{t}$ is the noise offset at time $t$; and
  • $(1-T_{t})$ represents the remaining ink portion for simulating absorption.

### v3

@@ Introduction
Wushu also serves as a virtual platform to review and appreciate the audience-generated works; the audience assumes the role of both writers and readers in the context of the “book.” Creating (writing) and appreciating (reading) the generated calligraphy becomes a cyclical process within this infinite “book.” Wushu can stimulate further attention and discussions concerning Chinese characters and calligraphy, such as their visual forms, linguistic meanings, and cultural essence.
@@
However, Wushu has its limitations, challenges, and opportunities that we would like to address in the future. The compendium is an essential component of Wushu’s concept and the archive. The audience finds it attractive that their collective efforts will form an actual “book.” However, we have not yet figured out the best format to present the archive, except for the real-time showcasing as part of Wushu’s interactive experiences. We currently save the digital pages as image files. Still, we are considering an online gallery that updates in real-time and publishes the archive publicly online or a physical book in print that mimics the reading experience. We will further explore and investigate how to present the archive better in the future.
During the exhibition of Wushu, we notice a significant difference in the performance of the English and Chinese caption models. The English caption model we acquired generates more accurate and fluent captions than the Chinese one. We attributed this difference not to the lack of AI research on Chinese captioning but to the scarcity and quality of the data sources. The large models that we used for captioning were trained with semantic information from various online data, while the predominant source of Chinese texts is of lower quality and diversity than the English data. Although this manuscript and research do not address the data and AI bias issues, we intend to explore them further.
Acknowledgments
Aven Le Zhou created the artwork discussed in this manuscript, encompassing its conceptual framework and technical development. He authored the major part of the manuscript unless specifically attributed otherwise. The artwork was shown and documented in various venues, such as the Ars Electronica Festival 2024, CVPR 2024 AI Art track and Digital Art China 2024. Jiayi Ye contributed insights into broader calligraphy practice and authored the calligraphy-related sections, including the Essential Principles in Calligraphy and parts of the Introduction. Tianchen Liu provided expertise in technical expressions for the Wushu: the Infinite Book. Kang Zhang guided the overall project and manuscript development as the PhD supervisor.